# Learning to Edit Visual Programs with Self-Supervision

### v1

@@ 1 Introduction
Though these prior approaches have made impressive progress, a common limitation is that they operate within the aforementioned one-shot paradigm. For instance, when using an autoregressive network, these systems will condition on a visual target and iteratively predict program tokens until completion. While this sequential inference procedure is sometimes wrapped in a more complex outer-search (e.g. beam-search or sequential Monte Carlo [10]), is allowed to reason over partial program executions [25], or is given access to executor-gradients that guide an inner-loop search [12, 44], all of these paradigms are distinct from how people typically write programs.
@@ 2 Related Work
However, most visual domains of interest lack such annotated data. As a result, a host of techniques have been investigated for this unsupervised setting, including: reinforcement learning [33, 10], differentiable execution architectures [44, 43, 20, 30, 35], learned proxy executors [37, 8], or bootstrapping methods [11, 19, 15]. All of these works operate within the aforementioned one-shot paradigm. Of note, SIRI investigates how analytical code rewriting operations can improve visual program inference networks in a bootstrapped learning paradigm [12]. Our system shares a similar motivation in that we aim to rewrite visual programs in a goal-directed fashion. However, instead of modifying programs with domain-specific fixed operations (e.g. differentiable parameter optimization), we explore a generalized alternative by introducing a network that learns how to edit programs.
@@
With the growing attention around the abilities of Large Language Models (LLMs), a number of recent works have explored how LLMs can be used to fix programs from input/output examples [34, 5, 23, 24] Though differing in details, the typical formulation these methods take involves presenting an LLM with a previous program version, and asking it to either (i) debug exceptions or (ii) modify program behavior in light of input/output mismatches. While these initial forays show promise, LLMs have not yet been able to write complex visual programs (in part due to poor visual reasoning capabilities), and even for more general program synthesis tasks the performance gains of code-editing LLMs are not definitive [26].
@@ 3 Method
In this section, we present our approach for learning how to edit visual programs. First we formalize our task of unsupervised visual program induction. For a particular domain, we are given a domain-specific language $L$ and an executor $E$ that converts programs $z$ from $L$ into visual outputs $x$. Given visual inputs from a target visual dataset that lacks program annotations, $x^{*}$in $X^{*}$, our goal is to find find $z^{*}$in $L$, such that $E$($z^{*}$$\sim$ $x^{*}$. This measure of similarity is usually checked under a domain specific reconstruction metric $M$.
A general approach employed by prior visual program induction works is to use an autoregressive model (e.g. a Transformer) conditioned on a visual encoding to predict a well-reconstructing program: $p(z|x)$. These one-shot models iteratively predict the next program token until the program is complete. We present a framework that employs a similar autoregressive model, but instead of predicting a complete program from scratch, we instead predict a local edit that modifies an input program. In the rest of this section, we first present how we design our edit network (Sec. 3.1). Then we discuss our unsupervised training procedure where we jointly finetune an edit network along with a one-shot network (Sec. 3.2. Finally, we describe how we combine these networks to search for visual programs (Sec. 3.3).
@@ 3.1 Edit Network Design
Edit Operations. There are many ways to parameterize the space of possible program edits. We choose to constrain the possible edit operations our network can produce by forcing it to select from a set of local editing operations designed for visual programs. For instance, for functional visual programming DSLs with transformation and combinator functions, we allow for six different edit operations: modifying a transform’s parameters (MP), modifying a transform (MT), adding a transform (AT), removing a transform (RT), removing a combinator (RC), or adding a combinator (AC). We provide more details in Appendix D. Some of these edit operations do not take in parameters (removing a transform) while others require new parameters (e.g. to modify the parameters of a transform we need to know the new parameters). Each of these edit operations can be applied to a program at a specific token location, and results in a local change. Subsequently, we task our edit network with predicting three items: an edit operation type, a location for that edit operation, and any extra parameters that operation requires.
@@
To predict the edit operation type, we take the output Transformer embedding from the first index of input program sequence. This embedding is sent through a linear layer which predicts a distribution over the possible edit operation types (yellow boxes, Fig. 1). To predict the edit operation location, we consider the embeddings that the Transformer produces over the tokens of the input program. Each of these location codes is sent through a linear layer, which predicts a value for each operation type. For a chosen operation type, we then normalize these values into a probability distribution across the length of the input program sequence (dark-blue boxes, Fig. 1). This distribution models the likelihood of where a specific edit operation type should be applied.
@@
Finally, we use our network to auto-regressively sample any extra parameters that a chosen edit operation might require. To accomplish this, we first slightly reformat the input program by inserting a special ‘sentinel token’ [28] associated with the chosen edit operation in two places: (1) at the specified edit operation location and (2) at the end location of the current program (\$AT, Fig. 1). This ‘sentinel’ tokens allows the network to know what operation is being applied to which position. Then, starting from the location of the second sentinel token, we can use the network to iteratively generate a sequence of parameter predictions with causal attention-masking, until an ‘END’ token is chosen (green boxes, Fig. 1).
@@
Training. Given an input program, how do we know which edit operations are helpful? If we have access to a target program, we can find a set of edit operations that would transform the input program into this target. We follow this logic to source training data for our edit network: given a start program and an end program, we analytically identify a set of edit operations that would bring about this transformation with a findEdits step. We can then convert this set of edit operations into a large set of (input, output) pairs that our network can train on. We provide further details on this algorithm in Appendix D. Once we have sourced paired data, we can train our network in a supervised fashion with a cross-entropy loss on the predicted operation type, location, and each parameter token (through teacher-forcing). Though we lack known programs for the target domain of interest, we next discuss a bootstrapped finetuning procedure that can work-around this issue.
3.2 Learning Paradigm
PLAD Finetuning. We begin with an overview of the PLAD method, which is depicted with the black text in Alg. 1 (see [19] for details). At the start of each round, the program inference network $p(z|x)$ is run over the target dataset $X^{*}$; the results of this inference procedure populate the entries of a best programs data-structure $P^{\text{BEST}}$ according to $M$. Then an unconditional generative model $p(z)$ is trained over the entries of $P^{\text{BEST}}$, and a set of ‘dreamed programs’, $P^{G}$, are sampled from this network. The weights of $p(z|x)$ are then finetuned using paired data sourced from $P^{\text{BEST}}$ and $P^{G}$. These steps are repeated for a set number of rounds, or until convergence.
Edit Model Finetuning. The blue-colored lines in Alg. 1 indicate the modifications we make to the PLAD algorithm to incorporate our edit network. Lines 8-10 explain the training logic. First we use  $p(z|x)$ to sample a set of programs $P^{S}$ conditioned on the executed outputs of the generated programs $P^{G}$. Treating $P^{S}$ as the starting points and $P^{G}$ as the end points, we can then use our findEdits operation to find sets of edit operations $ES$ that would realize these transformations. This provides us with paired data that we can use to finetune the weights of the edit network as explained in the prior section.
Synthetic Pretraining. PLAD finetuning is typically initialized with a synthetic pretraining phase (L1, Alg. 1). During pretraining, random scenes are sampled from $L$, and $p(z|x)$ can be trained on the paired data produced by executing these samples. Similarly, as we discuss in the results section, we find it useful to ‘pretrain’ the edit network on synthetic data (L2, Alg. 1). While multiple formulations are possible here, we re-use the same logic shown on lines 8-10, except we replace the set of target programs $P^{G}$ with random programs sampled from $L$.
Algorithm 1 Finetuning
@@ 3.2 Learning Paradigm
3.3 Inference
With the above procedure we can train our edit network, but how can we use it to find improved visual programs? This question is not only relevant at test-time, but impacts bootstrapped training, as we use an inference operation to populate the entries of $P^{\text{BEST}}$(L:5, Alg. 1). As depicted in Fig. 2, we design a search procedure that combines the strengths of the one-shot and editing paradigms. This search procedure maintains a population of programs, which are evolved over a number of rounds. The initial population is produced by sampling $p(z|x)$ Then for each round, we use the edit network to sample sets of edits for every program in the current population. We apply each of these sampled edits, and then re-sample the population for the next round according to a ranking based on $M$.
This formulation has a number of advantages. Instead of starting from scratch, or with random samples, we allow $p(z|x)$ to produce initial rough program estimates. These guesses are then refined through mutations over a series of editing rounds that are all directed at improving similarity towards the visual target. In Section 4.5 we compare this algorithm against alternative formulations. Critically, by applying this joint inference procedure during finetuning we form a virtuous cycle: improving the inference strategy leads to better $P^{\text{BEST}}$ entries, which results in better training data for $p(z|x)$ and $p(e|z,x)$, which in turn allows us to find to better $P^{\text{BEST}}$ entries in subsequent finetuning rounds. Finally, we note that this formulation maintains a nice symmetry between $p(z|x)$ and $p(e|z,x)$: in out joint finetuning algorithm  $p(e|z,x)$ trains on sequences sourced from sampling $p(z|x)$, and in this way its training distribution of edit operations well matches the population used to initialize the inference algorithm.
@@ 4 Results
We evaluate our edit network with experiments over multiple domains. First we describe our experimental design (Sec. 4.1). Then we compare the ability of different methods to accurately infer visual programs in terms of reconstruction performance (Sec. 4.2). We analyze how this performance changes as a function of time spent on inference (Sec. 4.3) or the size of the training target dataset (Sec. 4.4). Finally, we discuss results of our ablation study in Section 4.5.
@@ 4.1 Experimental Design
We provide a high-level overview of our experimental design. See Appendix C for details.
Methods. We compare our approach (OS+Edit) against the alternative of using only a one-shot model (OS Only). As described in Section 3 our approach jointly finetunes an edit network along with a one-shot network, and uses both of these networks to infer visual programs (Fig. 2). To control for the added time cost incurred by our inference procedure, we adapt a sampling-based inference loop for the OS Only variant, which we find results in a surprisingly strong baseline.
Domains. We consider three VPI domains (See App B): Layout, 2D CSG, and 3D CSG. In the Layout domain, scenes are created by placing colored 2D primitives on a canvas, and optionally modifying them by changing their size, location, or forming a symmetry group. In constructive solid geometry (CSG), complex shapes are formed by combining simple shapes with boolean set operations (union, intersection, difference). Our 2D CSG and 3D CSG domains differ in terms of their primitive types (e.g. squares vs cuboids) and the parameterizations of transformation functions (generalizing notions of scaling, translating, rotating, and symmetry grouping from $\mathds{R}^{2}$ to $\mathds{R}^{3}$).
Network Details. For each domain, we implement $p(z|x)$ as a decoder-only Transformer [38] that conditions on a set of visual tokens and predicts up to a maximum sequence length SL. Similarly, we implement $p(e|z,x)$ as a Transformer with the same architecture, except that it conditions on both (i) two sets of visual tokens and (ii) an input program of length SL, and it is only allowed to predict edit parameters up to a length of EL. Our visual encoders are all standard CNNs. For Layout we use a 2D-CNN that takes in an RGB 64x64 image, for 2D CSG we use a 2D-CNN that takes in a binary 64x64 image, and for 3D CSG we use a 3D-CNN that takes in a $32^{3}$ voxel field.
@@
Target Data. Like prior bootstrapping methods, our finetuning algorithm specializes our networks towards a target dataset of interest ($X^{*}$) that lacks known programs. For 2D CSG we use shapes from the dataset introduced by CSGNet [33], originally sourced from Trimble 3D warehouse. For 3D CSG we use shapes from the dataset introduced by PLAD [19], originally sourced from ShapeNet [4]. While we use the same test-sets as prior work (3000 / 1000 for CSG 2D / CSG 3D), we find that our method is able to offer good performance with much less training data. In our base experiments, we use 1000/100 train/val shapes for 2D CSG (from 10000 / 3000 available) and and 1000/100 train/val shapes for 3D CSG (from 10000 / 1000 available). For the Layout domain, we use the manually designed scenes sourced from [18] (1000 train / 100 val / 144 test).
@@ 4.2 Reconstruction Accuracy
We compare our OS+Edit approach against OS Only on each method’s ability to infer visual programs that accurately reconstruct test-set inputs in Table 1. As demonstrated, our joint finetuning paradigm that combines an edit network with a one-shot network consistently improves reconstruction performance. In these experiments, we ensure that each method gets to spend the same amount of time on inference by setting search parameters so that the average inference time per shape was equal: $\sim$ 5, $\sim$ 10, $\sim$ 60 seconds per shape for Layout, 2D CSG, and 3D CSG respectively. For OS Only, we use a sampling-based inference search where the model samples a population of complete programs for a set number of rounds. Though this approach was not as effective as combining our edit networks with one-shot initialization, it provides a surprisingly strong baseline. For instance, its 2D CSG CD score of 0.156 offers improvement over a number of related methods, even some that assume access to executor-gradients [35, 20, 12, 19]. We visualize reconstructions from this experiment in Figure 3, and find that qualitative evidence supports the quantitative trends.

### v2

@@ 1 Introduction
Though these prior approaches have made impressive progress, a common limitation is that they operate within the aforementioned one-shot paradigm. For instance, when using an autoregressive network these systems will condition on a visual target and iteratively predict program tokens until completion. While this sequential inference procedure is sometimes wrapped in a more complex outer-search (e.g. beam-search or sequential Monte Carlo [10]), is allowed to reason over partial program executions [25], or is given access to executor-gradients that guide an inner-loop search [12, 45], all of these paradigms are distinct from how people typically write programs.
@@
We release code for our experiments at: https://github.com/rkjones4/VPI-Edit
@@ 2 Related Work
However, most visual domains of interest lack such annotated data. As a result, a host of techniques have been investigated for this unsupervised setting, including: reinforcement learning [34, 10], differentiable execution architectures [45, 44, 20, 31, 36], learned proxy executors [38, 8], or bootstrapping methods [11, 19, 15]. All of these works operate within the aforementioned one-shot paradigm. Of note, SIRI investigates how analytical code rewriting operations can improve VPI networks in a bootstrapped learning paradigm [12]. Our system shares a similar motivation in that we aim to rewrite visual programs in a goal-directed fashion. However, instead of modifying programs with domain-specific fixed operations (e.g. differentiable parameter optimization), we explore a generalized alternative by introducing a network that learns how to edit programs.
@@
With the growing attention around the abilities of Large Language Models (LLMs), a number of recent works have explored how LLMs can be used to fix programs from input/output examples [35, 5, 23, 24]. Though differing in details, the typical formulation these methods take involves presenting an LLM with a previous program version, and asking it to either (i) debug exceptions or (ii) modify program behavior in light of input/output mismatches. While these initial forays show promise, LLMs have not yet been able to write complex visual programs (in part due to poor visual reasoning capabilities), and even for more general program synthesis tasks the performance gains of code-editing LLMs are not definitive [26].
@@ 3 Method
In this section, we present our approach for learning how to edit visual programs. First we formalize our task of unsupervised visual program induction. For a particular domain, we are given a domain-specific language (DSL) $L$ and an executor $E$ that converts programs $z$ from $L$ into visual outputs $x$. Given visual inputs from a target visual dataset that lacks program annotations, $x^{*}$in $X^{*}$, our goal is to find find $z^{*}$in $L$, such that $E$($z^{*}$$\sim$ $x^{*}$. This measure of similarity is usually checked under a domain specific reconstruction metric $M$.
A general approach employed by prior visual program induction works is to use an autoregressive model (e.g. a Transformer) that is conditioned on a visual encoding to predict a well-reconstructing program: $p(z|x)$. These one-shot models iteratively predict the next program token until the program is complete. We present a framework that employs a similar autoregressive model, but instead of predicting a complete program from scratch, we instead predict a local edit that modifies an input program. In the rest of this section, we first present how we design our edit network (Sec. 3.1). Then we discuss our unsupervised training procedure where we jointly finetune an edit network along with a one-shot network (Sec. 3.2. Finally, we describe how we combine these networks to search for visual programs (Sec. 3.3).
@@ 3.1 Edit Network Design
Edit Operations. There are many ways to parameterize the space of possible program edits. We choose to constrain the possible edit operations our network can produce by forcing it to select from a set of local editing operations designed for visual programs. For instance, for functional visual programming DSLs with transformation and combinator functions, we allow for seven different edit operations: modifying a transform’s parameters (MP), modifying a transform (MT), adding a transform (AT), removing a transform (RT), modifying a combinator (MC), removing a combinator (RC), or adding a combinator (AC). We provide more details in Appendix E. Some of these edit operations do not take in parameters (removing a transform) while others require new parameters (e.g. to modify the parameters of a transform we need to know the new parameters). Each of these edit operations can be applied to a program at a specific token location, and results in a local change. Subsequently, we task our edit network with predicting three items: an edit operation type, a location for that edit operation, and any extra parameters that operation requires.
@@
To predict the edit operation type, we take the output Transformer embedding from the first index of input program sequence. This embedding is sent through a linear layer which predicts a distribution over the possible edit operation types (yellow boxes, Fig. 1).
To predict the edit operation location, we consider the embeddings that the Transformer produces over the tokens of the input program. Each of these location codes is sent through a linear layer, which predicts a value for each operation type. For a chosen operation type, we then normalize these values into a probability distribution across the length of the input program sequence (dark-blue boxes, Fig. 1). This distribution models the likelihood of where a specific edit operation type should be applied.
@@
Finally, we use our network to autoregressively sample any extra parameters that a chosen edit operation might require. To accomplish this, we first slightly reformat the input program by inserting a special ‘sentinel token’ [29] associated with the chosen edit operation in two places: (1) at the specified edit operation location and (2) at the end location of the current program (\$AT, Fig. 1). This ‘sentinel’ tokens allows the network to know what operation is being applied to which position. Then, starting from the location of the second sentinel token, we can use the network to iteratively generate a sequence of parameter predictions with causal attention-masking, until an ‘END’ token is chosen (green boxes, Fig. 1).
@@
Training. Given an input program, how do we know which edit operations are helpful? If we have access to not only a visual target, but also its corresponding program, we can find a set of edit operations that would transform the input program into this target. We follow this logic to source training data for our edit network: given a start program and an end program, we analytically identify a set of edit operations that would bring about this transformation with a findEdits function. We can then convert this set of edit operations into a large set of (input, output) pairs that our network can train on. We provide further details on this algorithm in Appendix E. Once we have sourced paired data, through teacher-forcing we can train our network in a supervised fashion with a cross-entropy loss on the predicted operation type, location, and each parameter token. Though we lack known programs for the target domain of interest, we next discuss a bootstrapped finetuning procedure that provides a work-around for this issue.
Algorithm 1 Network Training
@@
3.2 Learning Paradigm
PLAD Finetuning. We begin with an overview of the PLAD method, which is depicted with the black text in Alg. 1 (see [19] for details). At the start of each round, the program inference network $p(z|x)$ is run over the target dataset $X^{*}$; the results of this inference procedure populate the entries of a best programs data-structure $P^{\text{BEST}}$ according to $M$. Then an unconditional generative model $p(z)$ is trained over the entries of $P^{\text{BEST}}$, and a set of ‘dreamed’ programs, $P^{G}$, are sampled from this network. The weights of $p(z|x)$ are then finetuned using paired data sourced from $P^{\text{BEST}}$ and $P^{G}$. These steps are repeated for a set number of rounds, or until convergence.
Edit Model Finetuning. The blue-colored lines in Alg. 1 indicate the modifications we make to the PLAD algorithm to incorporate our edit network. Lines 8-10 explain the training logic. First we use  $p(z|x)$ to sample a set of programs $P^{S}$ conditioned on the executed outputs of the generated programs $P^{G}$. Treating $P^{S}$ as the starting points and $P^{G}$ as the end points, we can then use our findEdits operation to find sets of edit operations $ES$ that would realize these transformations. This provides us with paired data that we can use to finetune the weights of the edit network through teacher forcing, as explained in the prior section.
Synthetic Pretraining. PLAD finetuning is typically initialized with a synthetic pretraining phase (Alg. 1, line 1). During pretraining, random programs are sampled from $L$, and $p(z|x)$ can be trained on the paired data produced by executing these samples. Similarly, as we discuss in the results section, we find it useful to ‘pretrain’ the edit network on synthetic data (Alg. 1, line 2). While multiple formulations are possible here, we re-use the same logic shown on lines 8-10, except we replace the set of target programs $P^{G}$ with random programs sampled from $L$.
3.3 Inference Algorithm
With the above procedure we can train our edit network, but how can we use this network to find improved visual programs? This question is not only relevant at test-time, but also impacts bootstrapped training, as we run an inner-loop search to populate the entries of $P^{\text{BEST}}$(Alg. 1, line 5). As depicted on the right side of Figure 2, we design a search procedure that combines the strengths of the one-shot and editing paradigms. This search procedure maintains a population of programs, which are evolved over a number of rounds. The initial population is produced by sampling $p(z|x)$ Then for each round, we use the edit network to sample sets of edits for every program in the current population. We apply each of these sampled edits, and then re-sample the population for the next round according to a ranking based on $M$.
This formulation has a number of advantages. Instead of starting from a blank canvas, or with random samples, we allow $p(z|x)$ to produce initial rough program estimates. These guesses are then refined through mutations over a series of editing rounds that are all directed at improving similarity towards the visual target. In Section 4.5 we compare this algorithm against alternative formulations. Critically, by applying this joint inference procedure during finetuning we form a virtuous cycle: improving the inference strategy leads to better $P^{\text{BEST}}$ entries, which results in better training data for $p(z|x)$ and $p(e|z,x)$, which in turn allows us to find to better $P^{\text{BEST}}$ entries in subsequent finetuning rounds. Finally, we note that this formulation maintains a nice symmetry between $p(z|x)$ and $p(e|z,x)$: in out joint finetuning algorithm  $p(e|z,x)$ trains on sequences sourced from sampling $p(z|x)$, and in this way its training distribution of edit operations well matches the population used to initialize the inference algorithm.
@@ 4 Results
We evaluate our edit network with experiments over multiple domains. First we describe our experimental design (Sec. 4.1). Then we compare the ability of different methods to accurately infer visual programs in terms of reconstruction performance (Sec. 4.2). We analyze how this performance changes as a function of time spent on inference (Sec. 4.3) or the size of the training target dataset (Sec. 4.4). Finally, we discuss results of an ablation study on our method in Section 4.5.
@@ 4.1 Experimental Design
We provide a high-level overview of our experimental design. See Appendix D for details.
Methods. We compare our approach (OS+Edit) against the alternative of using only a one-shot model (OS Only). As described in Section 3, our approach jointly finetunes an edit network along with a one-shot network, and uses both of these networks to infer visual programs (Fig. 2). To control for the added time cost incurred by our inference procedure, we adapt a sampling-based inference loop for the OS Only variant, which we find results in a surprisingly strong baseline.
Domains. We consider three VPI domains (see App C): Layout, 2D CSG, and 3D CSG. In the Layout domain, scenes are created by placing colored 2D primitives on a canvas, and optionally modifying them by changing their size, location, or forming a symmetry group. In constructive solid geometry (CSG), complex shapes are formed by combining simple shapes with boolean set operations (union, intersection, difference). Our 2D CSG and 3D CSG domains differ in terms of their primitive types (e.g. squares vs cuboids) and the parameterizations of transformation functions: generalizing notions of scaling, translating, rotating, and symmetry grouping from $\mathds{R}^{2}$ to $\mathds{R}^{3}$.
Network Details. For each domain, we implement $p(z|x)$ as a decoder-only Transformer [39] that conditions on a set of visual tokens and predicts up to a maximum sequence length SL. Similarly, we implement $p(e|z,x)$ as a Transformer with the same architecture, except that it conditions on (i) two sets of visual tokens and (ii) an input program of length SL, and it is only allowed to predict edit parameters up to a length of EL. Our visual encoders are all standard CNNs. For Layout we use a 2D CNN that takes in an RGB 64x64 image, for 2D CSG we use a 2D CNN that takes in a binary 64x64 image, and for 3D CSG we use a 3D CNN that takes in a $32^{3}$ voxel grid.
@@
Target Data. Like prior bootstrapping methods, our finetuning algorithm specializes our networks towards a target dataset of interest, $X^{*}$, that lacks known programs. For 2D CSG we use shapes from the dataset introduced by CSGNet [34], originally sourced from Trimble 3D warehouse. For 3D CSG we use shapes from the dataset introduced by PLAD [19], originally sourced from ShapeNet [4]. While we use the same test-sets as prior work (3000 / 1000 for 2D CSG / 3D CSG), we find that our method is able to offer good performance with much less training data. In our base experiments, we use 1000/100 train/val shapes for 2D CSG (from 10000 / 3000 available) and and 1000/100 train/val shapes for 3D CSG (from 10000 / 1000 available). For the Layout domain, we use the manually designed scenes sourced from [18] (1000 train / 100 val / 144 test).
@@ 4.2 Reconstruction Accuracy
We compare our OS+Edit approach against OS Only on each method’s ability to infer visual programs that accurately reconstruct test-set inputs in Table 1. As demonstrated, our joint finetuning paradigm that combines an edit network with a one-shot network consistently improves reconstruction performance. In these experiments, we ensure that each method gets to spend the same amount of time on inference by setting search parameters so that the average inference time per shape was equal: $\sim$ 5, $\sim$ 10, $\sim$ 60 seconds per shape for Layout, 2D CSG, and 3D CSG respectively. For OS Only, we use a sampling-based inference search where the model samples a population of complete programs for a set number of rounds. Though this approach provides a strong baseline, it was not as effective as combining our edit networks with one-shot initializations. In fact, for the 2D CSG domain, our formulation achieves reconstruction scores that surpass the performance of related methods that assume access to executor-gradients. On the 2D CSG test-set, we achieve a Chamfer distance (CD) of 0.111 (lower is better), whereas UCSG-Net [20] gets a CD of 0.320, SIRI [12] gets a CD of 0.260, and ROAP [36] gets a CD of 0.210 . Note that as the DSL, architecture, objective, and inference procedures differ across these various works, it’s hard to make any absolute claims from this direct comparison. Nevertheless we would like to emphasize that our method’s reconstruction performance on this task is very strong in the context of the related literature. We visualize reconstructions from this experiment in Figure 3, and find that qualitative evidence supports the quantitative trends.